# Pressure, Reference Points, and Risk-Taking Behavior: Evidence from Bench Press Competitions

### v1

@@ Abstract
This study explores the impact of pressure and reference points on risk-taking behavior and success outcomes in competitive settings. Using a unique panel dataset from official bench press competitions, we decompose the effects of pressure from rivals on lifters’ risk-taking decisions and realized outcomes. Our findings reveal that pressure from lower- and higher-ranked rivals drives more aggressive weight attempts, particularly in earlier stages, with the likelihood of success increasing as pressure intensifies. In third and final attempts, pressure from higher-ranked competitors leads to more conservative choices. Counterfactual analysis shows that in the absence of pressure, lifters adopt more conservative strategies in second attempts and take greater risks in third attempts, though with reduced success probabilities. These results underscore the importance of strategic focus on personal goals and the potential benefits of disregarding external pressures in the later stages of competition.
Keywords: reference point, pressure, risk, benchpress
@@ 1 Introduction
Understanding the behavioral effects of reference points, particularly in the context of pressure and risk-taking, is a critical area of research. The core concept is that an individual’s assessment of an outcome is influenced not only by the outcome itself but also by how it compares to a reference point (Tversky and Kahneman 1992). A significant challenge in this field is disentangling the effects of pressure from a reference point on risk-taking behavior and realized outcomes. Unlike controlled experimental settings as in Schwerter (2024), where risk-taking can be modeled as a simple binary lottery, field data often presents complexities that make such decompositions difficult. In this paper, we utilize a unique sequential competition setting and extensive panel dataset from official bench press competitions, offering a rare opportunity to explore the interplay between pressure and risk-taking behavior. This dataset enables us to observe weight attempts, lifting outcomes, rankings, and the exogenous pressure exerted by rivals over time. By examining how pressure from reference points influences subsequent risk-taking decisions, we aim to enhance the understanding of behavior under uncertainty and contribute to discussions on optimizing competition design to maximize each participant’s performance.
@@
The primary contribution of this study is the decomposition of the effects of pressure arising from reference points into distinct elements: the choice of weight attempt (analogous to a lottery choice) and the actual probability of success (analogous to a lottery gain). We also explore alternative competition designs that manipulate the availability of pressure-related information. The structure of official bench press competitions provides an ideal context to study the relationship between reference points and risk-taking behavior within the sequential game setting of weight attempts and weight lifting stages, where exogenous pressure is present. Specifically, a weight attempt under pressure from rivals can be seen as a risk-taking behavior similar to making a lottery choice. Whether the lift is successful depends on the pressure conditional on the weight attempt and the outcomes and attempts of rivals. Through this investigation, we seek to answer an analogous question in real-world competitive settings: What would be the impact on risk-taking behavior and the distribution of expected outcomes if the pressure exerted by rivals were removed?
Our results demonstrate pressure’s significant role in influencing lifters’ behavior during competition. We find that male lifters, as well as those with higher body weight and experience, are more likely to attempt higher weights in both their second and third attempts out of the three allowed. Notably, pressure from both lower and higher-ranked rivals during the second attempt leads to increased attempt weights, reflecting lifters’ tendency to challenge themselves under competitive conditions. However, in third and final attempts, the pressure from lower-ranked rivals continues to drive higher attempts, while pressure from higher-ranked rivals results in more conservative attempts, likely due to the lifter’s perceived unlikelihood of overtaking higher-ranked competitors. This highlights the influence of both competitive positioning and loss aversion on decision-making, with cautious behavior prevailing when lifters face strong upward competition in later stages.
In terms of success probability, lifters facing pressure during their second attempts exhibit a higher likelihood of success, driven by the fear of losing rank. However, by the third attempt, fatigue and the diminishing return on rank improvement cause pressure from lower-ranked rivals to have a negligible effect on success, while pressure to outperform higher-ranked rivals remains a motivating factor.
Simulations of counterfactual scenarios, where pressure factors are removed, show that lifters are more conservative in their second attempts and take on more risk in their third attempts, often attempting weights closer to or above their best. However, success probabilities decline under these conditions, especially for second attempts, as the absence of competitive pressure reduces the effort exerted by lifters. These findings suggest that while lifters can adjust their strategy in response to pressure, maintaining focus on committing to their personal goals and ignoring external pressure may be key to maximizing success, particularly in later attempts. In broader contexts such as maximizing expected gains from risky choices in similar sequential setting, individuals should commit to their original risky decisions, especially in the later stages of competition, ignoring external pressures from competitors.
@@ 1.1 Related literature
Second, this paper contributes to the literature on the effect of reference points on risk-taking actions. The most closely related study is Puente-Moncayo (2019), which examines the effect of current scores on the tendency to take high-risk shots in tennis. The author follows Ely et al. (2017), using the approach developed by Klaassen and Magnus (2009) (a structural model) to show that first serves should be riskier than second serves, and using two first serves (or two second serves) is suboptimal. In the experimental economics literature, Schwerter (2024) conducted experiments on risk-taking behavior when subjects were exposed to exogenously predetermined peer earnings. The author found that subjects exposed to higher peer earnings increased their risk-taking behavior. Similarly, Post et al. (2008) highlight the relevance of risk-taking behavior in path-dependent contexts, using data from the TV show “Deal or No Deal” and a replicated experiment where subjects choose to either “deal” for a sure prize or “no deal” to continue for an uncertain prize. Our paper is more ideal than these examples for several reasons. First, the weight selection conditional on personal bests and rivals’ outcomes is analogous to choosing a lottery, where the probability of successfully lifting the weight can be inferred from data by both lifters and econometricians. Second, lifting weights involves a simple, monoarticular action, which leads to relatively fewer external factors influencing the outcomes, compared to other sports.
@@ 2.1 Data source
Raw Male 175829 0.86 0.35 0.00 1.00
Personal best 175829 74.73 81.06 0.00 307.50
First attempt weight 175829 130.96 44.77 5.00 330.00
Second attempt weight 175829 137.42 45.89 20.00 352.50
Third attempt weight 175829 141.62 46.56 20.00 352.50
Successful first attempt 175829 0.86 0.35 0.00 1.00
Successful second attempt 175829 0.72 0.45 0.00 1.00
Successful third attempt 175829 0.40 0.49 0.00 1.00
Best attempt 168959 137.97 46.24 0.00 325.00
Age 175829 31.77 11.21 13.00 69.00
Bodyweight 175829 85.35 20.37 24.90 240.00
Single-ply Male 76605 0.80 0.40 0.00 1.00
Personal best 76605 75.12 92.05 0.00 402.50
First attempt weight 76605 156.17 64.67 5.00 445.00
Second attempt weight 76605 162.84 65.51 20.00 445.00
Third attempt weight 76605 168.06 66.65 20.00 445.00
Successful first attempt 76605 0.77 0.42 0.00 1.00
Successful second attempt 76605 0.65 0.48 0.00 1.00
Successful third attempt 76605 0.40 0.49 0.00 1.00
Best attempt 70687 160.23 64.68 0.00 445.00
Age 76605 32.32 12.93 13.00 69.00
Bodyweight 76605 84.90 23.19 30.50 245.00
Multi-ply Male 14628 0.97 0.18 0.00 1.00
Personal best 14628 79.40 107.58 0.00 425.00
First attempt weight 14628 197.02 63.56 25.00 480.00
Second attempt weight 14628 205.79 64.59 30.00 500.50
Third attempt weight 14628 212.03 65.67 30.00 520.00
Successful first attempt 14628 0.70 0.46 0.00 1.00
Successful second attempt 14628 0.57 0.49 0.00 1.00
Successful third attempt 14628 0.35 0.48 0.00 1.00
Best attempt 12820 201.35 63.09 0.00 430.00
Age 14628 32.17 10.08 13.00 69.00
Bodyweight 14628 94.65 19.64 34.40 207.80
@@
Table 2 presents summary statistics across three types of powerlifting equipment: raw, single-ply, and multi-ply. Each category includes attributes such as personal best, first, second, and third attempt weights, successful attempts, best attempt, age, and body weight. The raw category has the largest number of participants, with 175,829 individuals, an average personal best of 74.73 kg, and a mean bodyweight of 85.35 kg. The single-ply category includes 76,605 participants, with an average personal best of 75.12 kg and a mean bodyweight of 84.90 kg. In the multi-ply category, there are 14,628 participants, with an average personal best of 79.40 kg and a mean bodyweight of 94.65 kg.
Across all equipment types, the mean weights lifted increase from the first to the third attempts, and the percentage of successful attempts decreases with each successive attempt. This indicates that while lifters attempt heavier weights in successive attempts, the success rate declines. Approximately 80% of participants are male. The average age of participants is consistently around the early 30s across all categories. These statistics highlight key performance trends and demographic characteristics in competitive powerlifting, emphasizing differences in performance and success rates among different equipment types and attempts.
Personal best weights of lifters who have never participated in an official bench press competition are recorded as zero. In the empirical analysis, we focus on the second and third attempts and treat the realized outcome in the first attempt as the personal best for such lifters because the success probability in the first attempt is more than 90%.
@@ 2.2 Game setting
The objective of each lifter is to lift the maximum weight possible from a prone position on a flat bench to improve the own rank. In case of a tie (same weight lifted), the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) is ranked higher. Competitors are divided into categories based on weight class, age class, equipment class, and gender to ensure fair competition.
@@
Each lifter has three attempts.666 Lifting Procedure for each lifter at each attempt is divided into the following. 1. Setup: The lifter lies on their back on the bench, with feet flat on the floor or on the bench’s footrests, ensuring contact with the bench throughout the lift. 2. Grip: The lifter grips the barbell, usually slightly wider than shoulder-width apart. 3. Unracking: With the help of spotters if needed, the lifter unracks the barbell and holds it with arms fully extended above their chest. 4. Lowering: The barbell is lowered in a controlled manner to touch the chest, ensuring it is stable and paused for a moment. 5. Press Command: The judge gives a “Press” command once the barbell is motionless on the chest. The lifter then presses the barbell back to the starting position with arms fully extended. 6. Rack Command: Upon completion, and with the barbell in a stable position, the judge gives the “Rack” command, signaling the lifter to return the barbell to the rack. Each attempt is judged based on control, stability, and completion of lifting. A lift is considered successful if the lifter adheres to all the rules and completes the lift as per the judge’s commands from the viewpoints of two of three referees. After completing an attempt, the lifter must declare the weight for their next attempt within a specific time frame (usually 1 minute after their attempt). Once a weight is declared, it generally cannot be decreased, only increased. The order of lifters is adjusted after each round of attempts based on the declared weights for the next attempt, again proceeding from lightest to heaviest. This ensures that the competition is efficient and fair.
@@
To highlight the available information of each lifter at each attempt, consider a competition with three lifters (A, B, and C). The timeline of the first and second attempts is shown in Figure 1. For opening attempts, lifters $A,B$, and $C$ declare 100 kg, 105 kg, and 110 kg before the competition. Then, the order will be $A,B,C$. Lifter $A$ attempts $100$kg then the outcome is realized. Next, lifter $B$ attempts $105$kg. At the same time, lifter $A$ declares the next attempt weight $110$kg in one minute conditional on $A$’s own first outcome and $B$ and $C$’s attempt weights before $B$’s outcome is realized because lifter $B$ is trying 105kg at that time. Lifter $B$’s outcome is realized. Next, lifter $C$ attempts $110$kg. At the same time, lifter $B$ declares the next attempt weight $115$kg in one minute knowing $B$’s own first outcome, $A$’s outcome and next attempt, and $C$’s attempt weight before $C$’s outcome is realized because lifter $C$ is trying 110kg at that time. Lifter $C$’s outcome is realized. Lifter $C$ declares the next attempt weight $112.5$kg in one minute knowing $A$ and $B$s’ outcomes, their next attempts, and $C$’s own first outcome. For the second attempts, the new order will be A (110 kg), C (112.5 kg), B (115 kg). After the same procedure, the new order for the third attempts, will be similarly determined by the declared weights at the second attempts.

### v2

# Reference Points, Risk-Taking Behavior, and Competitive Outcomes in Sequential Settings

@@ Abstract
Understanding how competitive pressure affects risk-taking is crucial in sequential decision-making under uncertainty. This study examines these effects using bench press competition data, where individuals make risk-based choices under pressure. We estimate the impact of pressure on weight selection and success probability. Pressure from rivals increases attempted weights on average, but responses vary by gender, experience, and rivalry history. Counterfactual simulations show that removing pressure leads many lifters to select lower weights and achieve lower success rates, though some benefit. The results reveal substantial heterogeneity in how competition shapes both risk-taking and performance.
Keywords: Reference Point, Pressure, Risk, Sports Competition
@@ 1 Introduction
Understanding the behavioral effects of reference points, particularly in the context of pressure and risk-taking, is a critical area of research. The core concept is that an individual’s assessment of an outcome is influenced not only by the outcome itself but also by how it compares to a reference point (Tversky and Kahneman 1992). A significant challenge in this field is disentangling the effects of pressure from a reference point on risk-taking behavior and outcomes. Unlike controlled experimental settings as in Schwerter (2024), where risk-taking can be modeled as a binary lottery, field data presents complexities that make such decompositions difficult. In this paper, we utilize a sequential competition setting and panel dataset from official bench press competitions, offering an opportunity to explore the interplay between pressure and risk-taking behavior. This dataset enables us to observe weight attempts, outcomes, rankings, and the exogenous pressure exerted by rivals over time. By examining how pressure influences subsequent risk-taking decisions, we aim to enhance the understanding of behavior under uncertainty and contribute to discussions on optimizing competition design to maximize performance.
@@
The primary contribution of this study is the decomposition of the effects of pressure arising from reference points into distinct elements: the choice of weight attempt (analogous to a lottery choice) and the probability of success (analogous to a lottery gain). We also explore alternative competition designs that manipulate the availability of pressure-related information. The structure of official bench press competitions provides an ideal context to study the relationship between reference points and risk-taking within the sequential game setting of weight attempts and weight lifting stages, where exogenous pressure is present. Specifically, a weight attempt under pressure from rivals can be seen as a risk-taking behavior similar to making a lottery choice. Whether the lift is successful depends on the pressure conditional on the weight attempt and the rivals’ outcomes. Through this investigation, we seek to answer an analogous question in real-world competitive settings: What would be the impact on risk-taking behavior and expected outcomes if the pressure exerted by rivals were removed?
Our results demonstrate the role of pressure in shaping lifters’ behavior during competition. Pressure from both lower and higher-ranked rivals leads to increased attempt weights across both the second and third attempts, reflecting lifters’ tendency to challenge themselves under competition.
Regarding success probability, lifters under pressure from below show a slightly higher likelihood of success, possibly due to the motivation to defend their rank. However, when attempting to surpass higher-ranked rivals, success rates consistently decline, highlighting the difficulty of overtaking stronger competitors under pressure. This pattern is most pronounced in the third attempt, where fatigue and psychological strain may amplify the challenge. These results suggest that while competitive pressure can encourage bold strategies, it also introduces constraints that limit execution success in decisive moments.
Heterogeneity analyses further reveal that responses to pressure vary by gender, experience, and rivalry history. Male lifters exhibit stronger responses to both lower and higher rival pressure in attempt selection, whereas female lifters show a more muted response, especially under upward pressure. Female lifters experience a sharper decline in success probability when attempting to turn around a higher-ranked rival, suggesting a heightened sensitivity to stress. More experienced lifters are more sensitive to higher-ranked competitors, reinforcing the idea that competitive awareness strengthens with experience. Additionally, lifters with a history of repeated rivalry encounters show heightened sensitivity to pressure, suggesting that familiarity enhances their ability to navigate competition. For success probabilities, while pressure from lower-ranked rivals can improve outcomes, particularly among experienced lifters, upward pressure generally reduces them, with stronger effects among female competitors. Lifters with strong rivalry history show a small negative impact from upward pressure but a large positive impact from lower pressure. These confirm substantial heterogeneity.
To examine how pressure shapes the balance between ambition and feasibility in decision-making, we simulate counterfactual competition designs in which such pressure is removed. The results reveal consistent directional patterns—more conservative second attempts and more variable third attempts—but also considerable heterogeneity in both behavior and outcomes. While a subset of lifters benefit from the absence of pressure, many experience lower success rates and reduced expected outcomes, particularly in the second attempt. These differences reflect the uneven role of external pressure: for some, it enables focus and composure; for others, it disrupts execution or leads to overly cautious choices.
More broadly, in decision-making contexts involving sequential risky choices, individuals may experience diverging outcomes depending on how they respond to the presence or absence of external competition. Particularly in the later stages of competition, ignoring external pressure and focusing on personal strategy may enhance outcomes for some, while others rely on competitive cues to sustain performance. These findings reinforce the idea that responses to pressure are highly individual, suggesting that the effectiveness of competition design depends on how one reacts to competition.
@@ 1.1 Related literature
Second, this paper contributes to the literature on the effect of reference points on risk-taking actions. The most closely related studies are Puente-Moncayo (2019) and Genakos and Pagliero (2012). Puente-Moncayo (2019) examines the effect of current scores on the tendency to take high-risk shots in tennis. The author follows Ely et al. (2017), using the approach developed by Klaassen and Magnus (2009) (a structural model) to show that first serves should be riskier than second serves, and using two first serves (or two second serves) is suboptimal. Genakos and Pagliero (2012) study the effects of interim ranking in itself on weight application and success probability using the data of the powerlifting competition in the Olympic Games and World and European Championships between 1990 and 2006. In the experimental economics literature, Schwerter (2024) conducted experiments on risk-taking behavior when subjects were exposed to exogenously predetermined peer earnings. The author found that subjects exposed to higher peer earnings increased their risk-taking behavior. Similarly, Post et al. (2008) highlight the relevance of risk-taking behavior in path-dependent contexts, using data from the TV show “Deal or No Deal” and a replicated experiment where subjects choose to either “deal” for a sure prize or “no deal” to continue for an uncertain prize. Our paper is more ideal than these examples for several reasons. First, the weight selection conditional on personal bests and rivals’ outcomes is analogous to choosing a lottery, where the probability of successfully lifting the weight can be inferred from data by both lifters and econometricians. Second, lifting weights involves a simple, monoarticular action, which leads to relatively fewer external factors influencing the outcomes, compared to other sports including powerlifting (Genakos and Pagliero 2012). Third, unlike Genakos and Pagliero (2012), we exploit the sequential competition design which explicitly defines revealed information to each player. Fourth, we use a universe of official bench press competitions with a richer sample size and set of variables, which have been known recently as the most important factors to detect behavioral effects in the sports data (Teeselink et al. 2023).
@@
Following the seminal model by Lazear and Rosen (1981), tournament theory has been extended in various directions to better reflect behavioral patterns and decision-making processes observed in real-world settings. Two particularly relevant strands of this literature focus on incorporating risk-taking decisions instead of traditional effort choices and integrating reference-dependent preferences into contest frameworks. Hvide (2002), Hvide and Kristiansen (2003), and Taylor (2003) incorporate risk-taking behavior into tournament models, analyzing how participants’ strategic risk choices can be influenced by their relative positions and incentives within the tournament structure. Nieken and Sliwka (2010) show that the correlation of risk outcomes affects how both leading and trailing participants adjust their risk strategies. Gill and Stone (2010) and Dato et al. (2018) extend tournament models by incorporating reference-dependent preferences, highlighting how loss aversion and outcome expectations can shape competitive behavior and effort provision. These extensions are closely related to our study, as they emphasize the importance of risk-taking behavior and reference-dependent preferences in tournament settings.
@@ 2.1 Data source
Raw Male 175983 0.86 0.35 0.00 1.00
Personal best 175983 124.39 58.77 0.00 310.00
First attempt weight 175983 130.96 44.77 5.00 330.00
Second attempt weight 175983 137.42 45.89 20.00 352.50
Third attempt weight 175983 141.62 46.57 20.00 352.50
Successful first attempt 175983 0.86 0.35 0.00 1.00
Successful second attempt 175983 0.72 0.45 0.00 1.00
Successful third attempt 175983 0.40 0.49 0.00 1.00
Best attempt 169111 137.97 46.25 0.00 325.00
Age 175983 31.77 11.22 13.00 69.00
Bodyweight 175983 85.35 20.37 24.90 240.00
Single-ply Male 76681 0.80 0.40 0.00 1.00
Personal best 76681 133.97 80.06 0.00 402.50
First attempt weight 76681 156.16 64.67 5.00 445.00
Second attempt weight 76681 162.82 65.51 20.00 445.00
Third attempt weight 76681 168.05 66.65 20.00 445.00
Successful first attempt 76681 0.77 0.42 0.00 1.00
Successful second attempt 76681 0.65 0.48 0.00 1.00
Successful third attempt 76681 0.40 0.49 0.00 1.00
Best attempt 70756 160.21 64.68 0.00 445.00
Age 76681 32.32 12.93 13.00 69.00
Bodyweight 76681 84.89 23.19 30.50 245.00
Multi-ply Male 14678 0.97 0.18 0.00 1.00
Personal best 14678 157.23 95.59 0.00 425.00
First attempt weight 14678 197.10 63.60 25.00 480.00
Second attempt weight 14678 205.88 64.64 30.00 500.50
Third attempt weight 14678 212.12 65.72 30.00 520.00
Successful first attempt 14678 0.70 0.46 0.00 1.00
Successful second attempt 14678 0.57 0.49 0.00 1.00
Successful third attempt 14678 0.35 0.48 0.00 1.00
Best attempt 12864 201.43 63.12 0.00 430.00
Age 14678 32.20 10.10 13.00 69.00
Bodyweight 14678 94.69 19.65 34.40 207.80
@@
Table 2 presents summary statistics across three types of powerlifting equipment: raw, single-ply, and multi-ply. Each category includes attributes such as personal best, first, second, and third attempt weights, successful attempts, best attempt, age, and body weight. The raw category has the largest number of participants, with 175,983 individuals, an average personal best of 124.39 kg, and a mean bodyweight of 85.35 kg. The single-ply category includes 76,681 participants, with an average personal best of 133.97 kg and a mean bodyweight of 84.89 kg. In the multi-ply category, there are 14,678 participants, with an average personal best of 157.23 kg and a mean bodyweight of 94.69 kg.
Across all equipment types, the mean weights lifted increase from the first to the third attempts, and the percentage of successful attempts decreases with each successive attempt. This indicates that while lifters attempt heavier weights in successive attempts, the success rate declines. Approximately 86% of participants in the raw category, 80% in the single-ply category, and 97% in the multi-ply category are male. The average age of participants is around 31.77 years for raw lifters, 32.32 years for single-ply lifters, and 32.20 years for multi-ply lifters. These statistics highlight key performance trends and demographic characteristics in competitive powerlifting, emphasizing differences in performance and success rates among different equipment types and attempts.
Personal best weights of lifters who have never participated in an official bench press competition are recorded as zero. In the empirical analysis, we focus on the second and third attempts and treat the realized outcome in the first attempt as the personal best for such lifters because the success probability in the first attempt is 86% for raw lifters, 77% for single-ply lifters, and 70% for multi-ply lifters.
@@ 2.2 Game setting
The objective of each lifter is to lift the maximum weight possible from a prone position on a flat bench to improve their own rank. In case of a tie (same weight lifted), the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) is ranked higher. Competitors are divided into categories based on weight class, age class, equipment class, and gender to ensure fair competition.
@@
Each lifter has three attempts.666 Lifting Procedure for each lifter at each attempt is officially divided into the following. 1. Setup: The lifter lies on their back on the bench, with feet flat on the floor or on the bench’s footrests, ensuring contact with the bench throughout the lift. 2. Grip: The lifter grips the barbell, usually slightly wider than shoulder-width apart. 3. Unracking: With the help of spotters if needed, the lifter unracks the barbell and holds it with arms fully extended above their chest. 4. Lowering: The barbell is lowered in a controlled manner to touch the chest, ensuring it is stable and paused for a moment. 5. Press Command: The judge gives a “Press” command once the barbell is motionless on the chest. The lifter then presses the barbell back to the starting position with arms fully extended. 6. Rack Command: Upon completion, and with the barbell in a stable position, the judge gives the “Rack” command, signaling the lifter to return the barbell to the rack. Each attempt is judged based on control, stability, and completion of lifting. A lift is considered successful if the lifter adheres to all the rules and completes the lift as per the judge’s commands from the viewpoints of two of three referees. After completing an attempt, the lifter must declare the weight for their next attempt within a specific time frame (usually 1 minute after their attempt). Once a weight is declared, it generally cannot be decreased, only increased. The order of lifters is adjusted after each round of attempts based on the declared weights for the next attempt, again proceeding from lightest to heaviest. This ensures that the competition is efficient and fair.
@@
To highlight the available information of each lifter at each attempt, consider a competition with three lifters (A, B, and C). The timeline of the first and second attempts is shown in Figure 1. For opening attempts, lifters $A,B$, and $C$ declare 100 kg, 105 kg, and 110 kg before the competition. Then, the order will be $A,B$, and $C$. Lifter $A$ attempts 100 kg then the outcome is realized. Next, lifter $B$ attempts 105 kg. At the same time, lifter $A$ declares the next attempt weight 110 kg in one minute conditional on $A$’s own first outcome and $B$ and $C$’s attempt weights before $B$’s outcome is realized because lifter $B$ is trying 105 kg at that time. Lifter $B$’s outcome is realized. Next, lifter $C$ attempts 110 kg. At the same time, lifter $B$ declares the next attempt weight 115 kg in one minute knowing $B$’s own first outcome, $A$’s outcome and next attempt, and $C$’s attempt weight before $C$’s outcome is realized because lifter $C$ is trying 110 kg at that time. Lifter $C$’s outcome is realized. Lifter $C$ declares the next attempt weight 112.5 kg in one minute knowing $A$ and $B$s’ outcomes, their next attempts, and $C$’s own first outcome. For the second attempts, the new order will be A (110 kg), C (112.5 kg), B (115 kg). After the same procedure, the new order for the third attempts, will be similarly determined by the declared weights at the second attempts.